Bruce receives a weekly salart of $350. Hel also recieves a 6% commission on the total dollar volume of all sales he makes. What must his dollar volume be in a week, if he is to make a total of $710.?
Answer by checkley77(12569)

350+.06x=710
.06x=710-350
.06x=2360
x=360/.06
x=$6,000 is the dollar volume.
Proof:
350+.06*6,000=710
350+360=710
710=710